Private nonprofit four-year colleges charge, on average, $27,790 per year in tuition and fees. The standard deviation is $7,121. Assume the distribution is normal. Let X be the cost for a randomly selected college. Round all answers to 4 decimal places where possible.

a. What is the distribution of X? X ~ N(,)

b. Find the probability that a randomly selected Private nonprofit four-year college will cost less than 29,366 per year.

c. Find the 78th percentile for this distribution. $ (Round to the nearest dollar.)
Answer #1

Solution :

a.

X N ($27790 , $7121)

b.

P(x < 29366) = P[(x - ) / < (29366 - 27790) / 7121]

= P(z < 0.22)

= 0.5871

Probability = 0.5871

c.

Using standard normal table ,

P(Z < z) = 78%

P(Z < 0.77) = 0.78

z = 0.77

Using z-score formula,

x = z * +

x = 0.77 * 7121 + 27790 = 33273

$33273